OpenStudy (anonymous):

1/2r-1/6r+3=2+1/6r+1 step by step plz

OpenStudy (anonymous):

r/3+3 = r/6+3 Add (-r/6-3) to both sides: r/6 = 0 Divide both sides by 1/6: r = 0

OpenStudy (mathmate):

I have a couple of suggestions for you: 1. parentheses mean a lot in maths. Without parentheses, we do math in the order of multiplication and division BEFORE addition and subtraction. 1/2r-1/6r+3=2+1/6r+1 means (1/2r) - (1/6r) + 3 = 2 + (1/6r) + 1 It is not clear if that's what you really mean. 2. If you post 5 similar problems without waiting for the answer, it is clear that you just want the answer and not how to do it. That defeats the purpose of learning, and is detrimental to your cognitive development. Think about how you're going to handle exam questions when the time comes. Hint: show some effort on your part, give the problems a try. This also enables people trying to help you to understand where you encounter difficulties, and better help you. Good luck!